# PointAD+: Learning Hierarchical Representations for Zero-shot 3D Anomaly Detection

### v2

@@ 3.4 Pixel representation and point representation
[…]
$\tilde{\mathcal{K}}(x_{ij}^{3d},x_{il}^{3d})=\frac{\mathcal{K}(x_{ij}^{3d},x_{il}^{3d})}{\sum_{x_{in}^{3d}\in\mathcal{N}_{x_{ij}^{3d}}}\mathcal{K}(x_{ij}^{3d},x_{in}^{3d})},$ $\kern 4.25006pt\tilde{q_{ij}^{m}}\hskip-1.42271pt=\hskip-8.5359pt\sum_{q_{il}^{m}\in\mathcal{N}_{q_{ij}^{m}}}\hskip-7.11317pt\tilde{\mathcal{K}}(x_{ij}^{3d},x_{il}^{3d})\cdot q_{il}^{m},$
@@
$\hat{q_{ij}^{m}}=\alpha\tilde{q_{ij}^{m}}+(1-\alpha)$ $q_{ij}^{m}.$
@@ 3.5 Hierarchical representation learning
The key to ZS 3D anomaly detection lies in the model’s ability to capture generic anomaly semantics, enabling it to detect anomalies across various objects. This requires addressing two questions: how to effectively represent 3D abnormality and how to enable the model to recognize generalized 3D abnormality. First, 3D anomalies in our framework are characterized by two aspects: implicit 3D abnormality on the 2D renderings (rendering layer) and explicit 3D abnormality reflected on geometry (geometry layer). After this, we aim to adapt CLIP to 3D anomaly detection. Since CLIP was originally pre-trained to align object semantics, such alignment harms the generalization capacity of CLIP to recognize anomalies on various objects. Therefore, we propose a hierarchical representation learning, from both rendering and geometry layers, to globally and locally optimize hierarchical textual representations. Furthermore, we introduce cross-hierarchy alignment to facilitate mutual anomaly learning across layers. This enables PointAD+ to produce generic text embeddings for aligning global and local anomaly semantics.
@@ 3.5.1 Implicit point representation learning on the rendering layer (PointAD)
MIL-based 3D representation learning. To fully incorporate rendering anomaly semantics into PointAD+, we respectively devised two losses to regularize the rendering anomalies from 3D global anomaly and local anomaly semantics. First, we compute the cosine similarity between rendering textual representations $g^{r}_{c}$ and global pixel representations $f_{i}$ in each view. As point clouds are projected from different views, the resulting renderings in each view reflect specific parts of point clouds. We use view-wise MIL to integrate global implicit representations $p$ and then align global labels to capture the global semantics. Formally, the global 3D loss is defined as:
$L_{I-3d}^{global}=\frac{1}{N}\sum_{i}\textrm{CE}(\frac{1}{K}\textstyle\sum_{f_{i}^{(k)}\in p_{i}}P(g^{r}_{c},f_{i}^{(k)}),\max{(y^{3d}_{i})}),$
@@
$\textstyle S_{i(a)}^{3d}=\frac{1}{K}\sum_{k}Seg(g^{r}_{a},p_{i}^{m(k)}),\textstyle S_{i(n)}^{3d}=\frac{1}{K}\sum_{k}Seg(g^{r}_{n},p_{i}^{m(k)}).$
@@
$\hskip-5.0pt\textstyle S_{i(a)}^{3d}\hskip-0.50003pt=\hskip-0.50003pt\frac{1}{K}\sum_{k}\Big((R^{(k)})^{(-1)}\textrm{Up}(Seg(g^{r}_{a},f_{i}^{m(k)}))\otimes M_{i}^{(k)}\Big).$ (3)
@@
$L_{I-3d}^{local}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(S_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(S_{i(a)}^{3d},y_{i}^{3d})\Big).$
MTL-based pixel representation learning. We further improve PointAD+ and PointAD point understanding by capturing 2D glocal anomaly semantics into the object-agnostic text prompt template. We treat the anomaly recognition for one rendering from the point cloud as a task. Hence, we formulate the anomaly semantics learning for multiple 2D renderings as MTL. MTL-based 2D representation learning is divided into two parts for respective alignment to 2D global and local anomaly semantics. For 2D global semantics, we use CrossEntropy to quantify the discrepancy between the textual representations and each global Pixel representation $f_{i}$. Global MTL-based 2D representation learning $L_{I-2d}^{global}$ is defined as:
$\textstyle L_{I-2d}^{global}=\frac{1}{NK}\sum_{i,k}\textrm{CE}(P(g^{r}_{c},f_{i}^{(k)}),\max{(y_{i}^{(k)})}).$
Also, we focus on 2D abnormal regions to understand pixel-level anomalies. As the anomaly regions are typically smaller than normal regions, we employ Focal Loss to mitigate the class imbalance besides Dice Loss. Let $\oplus$ denote the concatenation operation. Local MTL-based 2D representation learning $L_{I-2d}^{local}$ is given as follows:
$\begin{split}\textstyle L_{I-2d}^{local}=\frac{1}{NK}\sum_{i,k}\textrm{Focal}(S_{i(n)}^{2d(k)}\oplus S_{i(a)}^{2d(k)},y_{i}^{(k)})+\\ \textrm{Dice}(S_{i(n)}^{2d(k)},I-y_{i}^{(k)})+\textrm{Dice}(S_{i(a)}^{2d(k)},y_{i}^{(k)}).\end{split}$ (4)
@@
$L^{r}_{hybrid}=L_{I-3d}^{global}+L_{I-3d}^{local}+L_{I-2d}^{global}+L_{I-2d}^{local}.$
@@ 3.5.2 Explicit point representation learning on the geometry layer
$\textstyle T_{i(a)}^{3d}=Seg(g^{g}_{a},\hat{q}^{m}_{i}),$ (5)
$\textstyle T_{i(n)}^{3d}=Seg(g^{g}_{n},\hat{q}^{m}_{i}).$
Here, to incorporate the abnormal spatial relations within point clouds, we use Dice Loss to align explicit local point representation $q_{i}^{m}$ with the 3D anomaly semantics:
$L_{E-3d}^{local}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(T_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(T_{i(a)}^{3d},y_{i}^{3d})\Big).$
@@ 3.5.3 Cross-hierarchy contrastive alignment
$L^{cross}\hskip-2.84544pt=\hskip-1.42271pt\textrm{CE}(\frac{cos(g^{g}_{a},g^{r}_{a})}{cos(g^{g}_{a},g^{r}_{a})+cos(g^{g}_{a},g^{r}_{n})})\hskip-2.84544pt+\hskip-2.84544pt\textrm{CE}(\frac{cos(g^{g}_{n},g^{r}_{n})}{cos(g^{g}_{n},g^{r}_{a})+cos(g^{g}_{a},g^{r}_{n})})$
$+\textrm{CE}(\frac{cos(g^{r}_{a},g^{g}_{a})}{cos(g^{r}_{a},g^{g}_{a})+cos(g^{r}_{a},g^{g}_{n})})\hskip-2.84544pt+\hskip-2.84544pt\textrm{CE}(\frac{cos(g^{r}_{n},g^{g}_{n})}{cos(g^{r}_{n},g^{g}_{a})+cos(g^{r}_{n},g^{g}_{n})})$
@@ 3.6.1 Training
$L_{hierarchical}=L^{r}_{hybrid}+L^{g}+L^{cross}.$ (6)
(a) Inference schematic for ZS 3D and M3D detection.
@@ 3.6.2 Inference
Since our model provides a unified framework to understand anomaly semantics from point and pixel, it can not only perform ZS 3D anomaly detection but also M3D anomaly detection in a plug-and-pay way. Next, we will introduce the inference process in detail:
@@
ZS M3D inference. When the RGB counterpart is available for testing, PointAD+ could directly integrate RGB information by feeding RGB images to 2D branch to derive pixel representations. In Figure LABEL:fig:_inference_2, we project these pixel representations back to 3D branch to respectively compute the RGB anomaly score map and anomaly score as $A^{m(rgb)}_{i}=P(g_{c},f_{i}^{(rgb)})$ and $A_{i}^{s(rgb)}=G_{\sigma}(S_{i(a)}^{3d(rgb)})$. The final multimodal anomaly score map and anomaly score are defined as $A_{i}^{m(mod)}=G_{\sigma}\big(0.5A_{i}^{m}+0.5A^{m(rgb)}_{i}\big)$ and $A_{i}^{s(mod)}=0.5(0.5A_{i}^{s(rgb)}+0.5A_{i}^{s})+0.5\max{(A_{i}^{m(mod)})}$. Also, the final multimodal anomaly score map and anomaly score for PointAD are defined as $B_{i}^{m(mod)}=G_{\sigma}\big(0.5B_{i}^{m}+0.5A^{m(rgb)}_{i}\big)$ and $B_{i}^{s(mod)}=0.5(0.5A_{i}^{s(rgb)}+0.5B_{i}^{s})+0.5\max{(B_{i}^{m(mod)})}$.
@@ 4.1 Experiment Setup
Figure 5: Visualization for PointAD and PointAD+ on MVTec3D-AD and Eyecandies in ZS 3D and M3D anomaly detection. Each row visualizes the anomaly score maps of objects from cross-category and cross-dataset scenarios using the model trained on "cookie" from MVTec3D-AD.
@@
Baselines. As ZS 3D and M3D anomaly detection remain underexplored areas in the field, we make a great effort to provide these comparisons. We apply the original CLIP to our framework for 3D detection, called CLIP + Rendering. Also, we reproduce SOTA 3D recognition works including PointCLIP V2 [54] and Cheraghian [55] and adapt them for ZS 3D anomaly detection. We compare the SOTA 2D anomaly detection approach AnomalyCLIP [16] by fine-tuning it on depth maps. All hyperparameters in these baselines are kept the same. We will present the detailed reproduction:
  • AnomalyCLIP (ICLR 2024) is a SOTA zero-shot 2D anomaly detection method. AnomalyCLIP introduces object-agnostic learning to capture generic anomaly semantics of images. We adapt AnomalyCLIP in 3D detection by fine-tuning AnomalyCLIP on depth images.
TABLE II: Performance comparison on ZS M3D anomaly detection in "one-vs-rest" setting.
@@
TABLE IV: Performance comparison on ZS M3D anomaly detection in cross-dataset setting.
@@ 4.2 Main Results
We train PointAD and PointAD+ on three objects on MVTec3D-AD, Eyecandies, and Real3D-AD. Over three runs, the averaged results on one-vs-rest and cross-dataset settings are reported. We use metric pairs (I-AUROC% and AP%) and (P-AUROC% and AUPRO%) to evaluate the global and local detection performance, respectively. The best and second-best results in ZS are highlighted in Red and Blue. G. and L. represent 3D global and local anomaly detection. M3D global and local detection are abbreviated as MG. and ML.
@@
ZS M3D anomaly detection. PointAD+ and PointAD allow the direct incorporation of RGB counterpart when it is available during the test. Table II indicates that PointAD+ and PointAD leverage additional RGB information in a plug-and-plug manner to boost their performance. PointAD gets the performance gain from 82.0% to 86.9% AUROC and from 94.2% to 96.1% AP for global semantics on MVTec3D-AD. Additionally, as for local semantics, the performance improves from 95.5% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. A large performance gain is also obtained on Eyecandies and Real3D-AD. PointAD+ can also benefit from the RGB information and break the performance achieved by PointAD, i.e., 98.2% vs. 97.2% P-AUROC and 93.1% vs. 90.2% AUPRO on MVTec3D-AD, and 79.2% vs. 77.7% I-AUROC and 81.7% vs. 80.4% AP on Eyecandies. As other methods can improve their performance in some metrics, they still suffer from performance degradation in other metrics due to inefficient integration of the two modalities. Instead, PointAD+ and PointAD achieve overall improvement across all metrics by incorporating explicit joint constraints on both point and pixel information. Figure 5 also illustrates the superior performance of PointAD+ on ZS M3D detection.
Cross-dataset ZS anomaly detection. Cross-dataset detection is conducted to further evaluate the zero-shot capacity of PointAD+ and PointAD, using one object as an auxiliary and testing objects with totally different semantics and scenes in another dataset. We compare all baselines that need fine-tuning in Table III and M3D from Table IV. PointAD demonstrates strong cross-dataset generalization performance on Eyecandies and Real3D-AD, with nearly no obvious performance decay compared to the one-vs-rest setting. The strong transfer ability highlights its robust generalization capabilities in detecting anomalies in objects with unseen semantics and backgrounds. Compared to PointAD, PointAD+ exhibits a performance drop on Real3D-AD, with P-AUROC decreasing from 76.0% to 72.5%, I-AUROC dropping from 83.5% to 74.3%, and AP declining from 84.7% to 78.2%. The main advantage of PointAD+ over PointAD is to introduce explicit point learning and reveal the point relations. Hence, its performance gain depends on the gap in geometry between the source and target point clouds. Despite this, PointAD+ generally achieves comparable or superior performance to PointAD. We would like to point out that developing a more generalized point encoder to model point relations could help mitigate this spatial gap, thereby further improving its ZS detection capacity. We leave it for future research.

### v3

@@ 3.5.1 Implicit point space
3.5.2 Explicit point space
Since we interpret 3D abnormalities using 2D abnormalities through a 2D encoder in the implicit point space, this approach fails to consider the intrinsic geometry of the point cloud. Therefore, we seek to model explicit point representations to better capture geometric information within point clouds.
Explicit point representation: Explicit point representation requires obtaining the specific point representation for modeling abnormal spatial positions, rather than set-like implicit point representation. Since point information is represented as its counterpart pixels in our framework, we resort to aggregating the above implicit point representation to obtain explicit point representation:
$q_{i}^{m}=\{q_{ij}^{m}|q_{ij}^{m}=\ell(p_{ij}^{m(1)},...,p_{ij}^{m(K)})\}_{j=1}^{n},$ (3)
where $\ell(\cdot)$ represents the aggregation strategy, i.e., take the mean or maximum of the pixel representations as point representations or employing a parameterized network. If a parameterized network is adopted, it would bring additional computational overhead. To ensure computational efficiency, we adopt the mean operation as the aggregation strategy.
G-aggregation The above representations lack the inherent geometry information and further reduce the gap between the 2D and 3D representations. This makes it challenging for the point representation to reflect the relative spatial relationships within point clouds comprehensively. Hence, we propose G-aggregation to inject the geometry information into point representation aggregation, enabling the point representations to be aware of geometry characteristics. In G-aggregation, the process begins with an initial aggregation, where the explicit point representation is obtained by aggregating the implicit point representation. To further refine the initial explicit point representation with geometry details, we search the k nearest neighbors of $q_{ij}^{m}$ within distance space $\mathcal{G}$, i.e., $\mathcal{N}_{x_{ij}^{3d}}=\{x_{il}^{3d}\mid x_{il}^{3d}\in\operatorname{TopK}(\{x_{i}^{3d}\}\setminus x_{ij}^{3d},\mathcal{G}(x_{ij}^{3d},x_{il}^{3d}))\}$. In this work, we adopt Euclidean distance as the primary metric for measuring relative distances. When the neighboring points are selected, the second aggregation step is performed to refine the point representations. This is achieved by weighting the point representations of their neighbors, using a Gaussian kernel $\mathcal{K}(x_{ij}^{3d},x_{il}^{3d})=\exp\left(-\|x_{ij}^{3d}-x_{il}^{3d}\|^{2}/2\sigma^{2}\right)$ to assess their contributions, followed by normalization.
$\tilde{\mathcal{K}}(x_{ij}^{3d},x_{il}^{3d})=\frac{\mathcal{K}(x_{ij}^{3d},x_{il}^{3d})}{\sum_{x_{in}^{3d}\in\mathcal{N}_{x_{ij}^{3d}}}\mathcal{K}(x_{ij}^{3d},x_{in}^{3d})},$ $\kern 4.25006pt\tilde{q_{ij}^{m}}\hskip-1.42271pt=\hskip-8.5359pt\sum_{q_{il}^{m}\in\mathcal{N}_{q_{ij}^{m}}}\hskip-7.11317pt\tilde{\mathcal{K}}(x_{ij}^{3d},x_{il}^{3d})\cdot q_{il}^{m},$ (4)
@@ 3.5.2 Explicit point space
${\color[rgb]{0,0,0}\hat{q_{ij}^{m}}=\alpha\tilde{q_{ij}^{m}}+(1-\alpha)q_{ij}^{m}.}$ (5)
@@ 3.6 Hierarchical representation learning
The key to ZS 3D anomaly detection lies in the model’s ability to capture generic anomaly semantics, enabling it to detect anomalies across various objects. The above section focuses on effectively representing 3D anomalies. Here, we propose hierarchical representation learning to align these representations on two layers: the rendering/bottom layer aligns the implicit point representation and pixel representation with the learnable rendering prompt in implicit point space; the geometric/top layer aligns the explicit point representations with the learnable geometric prompts in explicit point space. Furthermore, we introduce cross-hierarchy alignment to facilitate mutual anomaly learning across layers. This enables PointAD+ to produce generic text embeddings for generalized anomaly semantics.
@@ 3.6.1 Implicit point representation learning on the rendering layer (PointAD)
MIL-based implicit point representation learning. To fully incorporate rendering anomaly semantics into PointAD+, we respectively devised two losses to regularize the rendering anomalies from 3D global anomaly and local anomaly semantics. First, we compute the cosine similarity between rendering textual representations $g^{r}_{c}$ and global pixel representations $f_{i}$ in each view. As point clouds are projected from different views, the resulting renderings in each view reflect specific parts of point clouds. We use view-wise MIL to integrate global implicit representations $p$ and then align global labels to capture the global semantics. Formally, the global 3D loss is defined as:
$L_{I(3d)}^{g}=\frac{1}{N}\sum_{i}\textrm{CE}(\frac{1}{K}\textstyle\sum_{f_{i}^{(k)}\in p_{i}}P(g^{r}_{c},f_{i}^{(k)}),\max{(y^{3d}_{i})}).$ (6)
@@
$\textstyle S_{i(a)}^{3d}$ $=\frac{1}{K}\sum_{k}Seg(g^{r}_{a},p_{i}^{m(k)})$ (7)
$\textstyle S_{i(n)}^{3d}$ $=\frac{1}{K}\sum_{k}Seg(g^{r}_{n},p_{i}^{m(k)}).$ (8)
@@
$\hskip-8.5359pt\textstyle S_{i(a)}^{3d}\hskip-0.50003pt=\hskip-0.50003pt\frac{1}{K}\sum_{k}\Big((R^{(k)})^{(-1)}\textrm{Up}(Seg(g^{r}_{a},f_{i}^{m(k)}))\otimes M_{i}^{(k)}\Big).$ (10)
@@
$\hskip-2.84544pt\textstyle L_{I(3d)}^{l}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(S_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(S_{i(a)}^{3d},y_{i}^{3d})\Big).$ (11)
MTL-based pixel representation learning. We further improve PointAD+ and PointAD point understanding by capturing 2D glocal anomaly semantics into the object-agnostic text prompt template. We treat the anomaly recognition for one rendering from the point cloud as a task. Hence, we formulate the anomaly semantics learning for multiple 2D renderings as MTL. MTL-based 2D representation learning is divided into two parts for respective alignment to 2D global and local anomaly semantics. For 2D global semantics, we use CrossEntropy to quantify the discrepancy between the textual representations and each global Pixel representation $f_{i}$. Global MTL-based 2D representation learning $L_{I(2d)}^{g}$ is defined as:
$\textstyle L_{I(2d)}^{g}=\frac{1}{NK}\sum_{i,k}\textrm{CE}(P(g^{r}_{c},f_{i}^{(k)}),\max{(y_{i}^{(k)})}).$ (12)
Also, we focus on 2D abnormal regions to understand pixel-level anomalies. As the anomaly regions are typically smaller than normal regions, we employ Focal Loss to mitigate the class imbalance besides Dice Loss. Let $\oplus$ denote the concatenation operation. Local MTL-based 2D representation learning $L_{I(2d)}^{l}$ is given as follows:
$\begin{split}\textstyle L_{I(2d)}^{l}=\frac{1}{NK}\sum_{i,k}\textrm{Focal}(S_{i(n)}^{2d(k)}\oplus S_{i(a)}^{2d(k)},y_{i}^{(k)})+\\ \textrm{Dice}(S_{i(n)}^{2d(k)},I-y_{i}^{(k)})+\textrm{Dice}(S_{i(a)}^{2d(k)},y_{i}^{(k)}).\end{split}$ (13)
@@
${\color[rgb]{0,0,0}L^{r}_{hybrid}=L_{I(3d)}^{g}+L_{I(3d)}^{l}+L_{I(2d)}^{g}+L_{I(2d)}^{l}.}$ (14)
@@ 3.6.2 Explicit point representation learning on the geometry layer
${\color[rgb]{0,0,0}\textstyle T_{i(a)}^{3d}=Seg(g^{g}_{a},\hat{q}^{m}_{i}),}$ (15)
${\color[rgb]{0,0,0}\textstyle T_{i(n)}^{3d}=Seg(g^{g}_{n},\hat{q}^{m}_{i}).}$ (16)
Here, to incorporate the abnormal spatial relations within point clouds, we minimize $L_{E-3d}^{local}$ ($L^{g}$) to align explicit local point representation $q_{i}^{m}$ with the 3D anomaly semantics:
$\color[rgb]{0,0,0}\textstyle L_{E-3d}^{local}=\frac{1}{N}\sum_{i}\Big(\textrm{Dice}(T_{i(n)}^{3d},I-y_{i}^{3d})+\textrm{Dice}(T_{i(a)}^{3d},y_{i}^{3d})\Big).$} (17)
(a) Inference schematic for ZS 3D and multimodal 3D detection.
Figure 5: Visualization for PointAD and PointAD+ on MVTec3D-AD and Eyecandies in ZS 3D and multimodal 3D anomaly detection. Each row visualizes the anomaly score maps of objects from cross-category and cross-dataset scenarios using the model trained on "cookie" from MVTec3D-AD.
@@ 3.6.3 Cross-hierarchy contrastive alignment
$L^{c}=$ $-\log\frac{\exp(\cos(g^{g}_{a},g^{r}_{a}))}{\exp(\cos(g^{g}_{a},g^{r}_{a}))+\exp(\cos(g^{g}_{a},g^{r}_{n}))}$
$-\log\frac{\exp(\cos(g^{g}_{n},g^{r}_{n}))}{\exp(\cos(g^{g}_{n},g^{r}_{n}))+\exp(\cos(g^{g}_{n},g^{r}_{a}))}$ (18)
$-\log\frac{\exp(\cos(g^{r}_{a},g^{g}_{a}))}{\exp(\cos(g^{r}_{a},g^{g}_{a}))+\exp(\cos(g^{r}_{a},g^{g}_{n}))}$
$-\log\frac{\exp(\cos(g^{r}_{n},g^{g}_{n}))}{\exp(\cos(g^{r}_{n},g^{g}_{n}))+\exp(\cos(g^{r}_{n},g^{g}_{a}))}$
@@ 3.7.1 Training
$L_{hierarchical}=L^{r}_{hybrid}+L^{g}+L^{c}.$ (19)
@@ 3.7.2 Inference
Since our model provides a unified framework to understand anomaly semantics from point and pixel, it can not only perform ZS 3D anomaly detection but also multimodal 3D anomaly detection in a plug-and-pay way. Next, we will introduce the inference process in detail:
@@
ZS multimodal 3D inference. When the RGB counterpart is available for testing, PointAD+ could directly integrate RGB information by feeding RGB images to 2D branch to derive pixel representations. In Figure LABEL:fig:_inference_2, we project these pixel representations back to 3D branch to respectively compute the RGB anomaly score map and anomaly score as $A^{m(rgb)}_{i}=P(g_{c},f_{i}^{(rgb)})$ and $A_{i}^{s(rgb)}=G_{\sigma}(S_{i(a)}^{3d(rgb)})$. The final multimodal anomaly score map and anomaly score are defined as $A_{i}^{m(mod)}=G_{\sigma}\big(0.5A_{i}^{m}+0.5A^{m(rgb)}_{i}\big)$ and $A_{i}^{s(mod)}=0.5(0.5A_{i}^{s(rgb)}+0.5A_{i}^{s})+0.5\max{(A_{i}^{m(mod)})}$. Also, the final multimodal anomaly score map and anomaly score for PointAD are defined as $B_{i}^{m(mod)}=G_{\sigma}\big(0.5B_{i}^{m}+0.5A^{m(rgb)}_{i}\big)$ and $B_{i}^{s(mod)}=0.5(0.5A_{i}^{s(rgb)}+0.5B_{i}^{s})+0.5\max{(B_{i}^{m(mod)})}$.
@@ 4.1 Experiment Setup
TABLE II: Performance comparison on ZS multimodal 3D anomaly detection in "one-vs-rest" setting.
@@
Baselines. As ZS 3D and multimodal 3D anomaly detection remain underexplored areas in the field, we make a great effort to provide these comparisons. We apply the original CLIP to our framework for 3D detection, called CLIP + Rendering. Also, we reproduce SOTA 3D recognition works including PointCLIP V2 [54] and Cheraghian [55] and adapt them for ZS 3D anomaly detection. We compare the SOTA 2D anomaly detection approach AnomalyCLIP [16] by fine-tuning it on depth maps. All hyperparameters in these baselines are kept the same. We will present the detailed reproduction:
  • AnomalyCLIP (ICLR 2024) is a SOTA ZS 2D anomaly detection method. AnomalyCLIP introduces object-agnostic learning to capture generic anomaly semantics of images. We adapt AnomalyCLIP in 3D detection by fine-tuning AnomalyCLIP on depth images.
@@
TABLE IV: Performance on ZS multimodal 3D anomaly detection in cross-dataset setting.
@@ 4.2 Main Results
We train PointAD and PointAD+ on three objects on MVTec3D-AD, Eyecandies, and Real3D-AD. Over three runs, the averaged results on one-vs-rest and cross-dataset settings are reported. We use metric pairs (I-AUROC% and AP%) and (P-AUROC% and AUPRO%) to evaluate the global and local detection performance, respectively. The best and second-best results in ZS are highlighted in Red and Blue. G. and L. represent 3D global and local anomaly detection. multimodal 3D global and local detection are abbreviated as MG. and ML.
@@
ZS multimodal 3D anomaly detection. PointAD+ and PointAD allow the direct incorporation of RGB counterpart when it is available during the test. Table II indicates that PointAD+ and PointAD leverage additional RGB information in a plug-and-plug manner to boost their performance. PointAD gets the performance gain from 82.0% to 86.9% AUROC and from 94.2% to 96.1% AP for global semantics on MVTec3D-AD. Additionally, as for local semantics, the performance improves from 95.5% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. A large performance gain is also obtained on Eyecandies and Real3D-AD. PointAD+ can also benefit from the RGB information and break the performance achieved by PointAD, i.e., 98.2% vs. 97.2% P-AUROC and 93.1% vs. 90.2% AUPRO on MVTec3D-AD, and 79.2% vs. 77.7% I-AUROC and 81.7% vs. 80.4% AP on Eyecandies. As other methods can improve their performance in some metrics, they still suffer from performance degradation in other metrics due to inefficient integration of the two modalities. Instead, PointAD+ achieves overall improvement by incorporating explicit joint constraints on both point and pixel information.
Cross-dataset ZS anomaly detection. Cross-dataset detection is conducted to further evaluate the ZS capacity of PointAD+ and PointAD, using one object as an auxiliary and testing objects with totally different semantics and scenes in another dataset. We compare all baselines that need fine-tuning in Table III and multimodal 3D from Table IV. PointAD demonstrates strong cross-dataset generalization performance on Eyecandies and Real3D-AD, with nearly no obvious performance decay compared to the one-vs-rest setting. The strong transfer ability highlights its robust generalization capabilities in detecting anomalies in objects with unseen semantics and backgrounds. Compared to PointAD, PointAD+ exhibits a performance drop on Real3D-AD, with P-AUROC decreasing from 76.0% to 72.5%, I-AUROC dropping from 83.5% to 74.3%, and AP declining from 84.7% to 78.2%. The main advantage of PointAD+ over PointAD is to introduce explicit point learning and reveal the point relations. Hence, its performance gain depends on the gap in geometry between the source and target point clouds. Despite this, PointAD+ generally achieves comparable or superior performance to PointAD. We would like to point out that developing a more generalized point encoder to model point relations could help mitigate this spatial gap, thereby further improving its ZS detection capacity. We leave it for future research.